A business idea requires to spend $20,000 on training the staff this year and will save $1,000 a every year afterwards (because the staff is more productive).
a. Will the firm implement the idea if the interest rate is 4%? Explain.
b. Will the firm implement the idea if the interest rate is 8%? Explain.
c. At what interest rate is the firm indifferent between implementing the idea and abandoning it? Explain.
1
Expert's answer
2019-06-24T08:49:54-0400

a. If the interest rate is 4%, then the net present value is NPV = -20,000 + 1,000/0.04 = 5,000, so the firm will implement the idea.

b. If the interest rate is 8%, then the net present value is NPV = -20,000 + 1,000/0.08 = -7,500, so the firm will not implement the idea.

c. The firm is indifferent between implementing the idea and abandoning it if NPV = 0, so the interest rate is: IRR = 1,000/20,000×100% = 5%.